# FP4DiT: Towards Effective Floating Point Quantization for Diffusion Transformers

### v2

@@ 4.2 Main Results
In our experiment, we consider four baseline approaches: Q-Diffusion (Li et al., 2023), TFMQ-DM (Huang et al., 2024), ViDiT-Q (Zhao et al., 2024) and Q-DiT (Chen et al., 2025). Note that while Q-Diffusion and TFMQ-DM are originally designed for U-Nets, we modify these approaches to use the same online token-wise activation quantization as FP4DiT per Table 1, while ViDiT-Q and Q-DiT uses this mechanism by default. Further, note that ViDiT-Q uses mix-precision meaning some of their layers are not quantized to 4-bits. For the sake of convenience, we use their mix-precision as a W4 baseline to conduct our experiments.
@@
We generate $512\times 512$ resolution images using PixArt-$\alpha$ and $1024\times 1024$ for PixArt-$\Sigma$ and Hunyuan. For evaluation, we primarily consider the Human Preference Score v2 (HPSv2) (Wu et al., 2023a) benchmark and MS-COCO 2014 (Lin et al., 2014a) validation set. Specifically, HPSv2 considers four image categories: animation, concept-art, painting and photography, and estimates the human preference score of an image generated using a prompt with respect to one category. Each category contains 800 prompts, requiring 3.2k images to be generated to fully evaluate. The final HPSv2 score is the average estimated human preference across all four categories. Additionally, for COCO 2014, we measure the FID (Heusel et al., 2017) and CLIP (Hessel et al., 2021) score using prompts from the validation set.
@@
Table 2 compares our method with the stated baseline approaches at the W4A8 and W4A6 precision levels. FP4DiT consistently outperforms all other methods across three different base models at every precision level in terms of each performance metric. On HPSv2, FP4DiT achieves the best average performance across every model and bit-precision combination. Although other approaches may achieve slightly higher performance on an individual category, these gaps are small, while FP4DiT leads in terms of FID and CLIP on the COCO 2014 validation set.
@@
Next, we compare performance across activation bit-precision. For the PixArt-series DiT models, A6 precision can cause substantial performance degradation for some baselines, while FP4DiT still maintains high performance. This is specifically noteworthy on PixArt-$\Sigma$ as baseline approaches suffer a loss in terms of either FID and/or CLIP performance at A6 compared to A8, but this is not the case for FP4DiT. For the Hunyuan DiT model, A6 quantization remains highly challenging across the board, however, our method still achives the best results at this level. Overall though, Table 2 demonstrates the efficacy of FP4DiT in term of quantitative T2I compared to prominant baseline approaches.
@@
Next, Figure 7 provides qualitative image samples on the PixArt models. Note the higher quality of the images generated by FP4DiT at both the A8 and A6 levels. Specifically, the puppies have more realistic detail and the generated image more closely aligns with the W16A16 model. This is especially true for the PixArt-$\Sigma$ sample images, where the FP4DiT show detailed, but not blurry northern lights while maintaining detail on the snowy landscape in the foreground.
@@ 4.4 Hardware Cost Comparison
We develop a CUDA 12.8 kernel and measure hardware latency on an Nvidia 5080, a Blackwell GPU which supports low-bit FP formats. We consider the self-attention weight/activation dimensions found in PixArt-$\alpha$/$\Sigma$ when generating an $1024\times 1024$ image. Table 6 reports our findings. The result indicates that, despite the general expectation that floating-point computations are more demanding than integer operations, the latency results are nearly identical across all tested quantization precisions. Specifically, although A6 precision imposes some additional overhead compared to A8 (either for FP or INT), this is quite minor, leading to an overall speedup around 1.5x compared to W16A16, much like the popular W4A8 integer quantization (Li et al., 2023; Zhao et al., 2024; Chen et al., 2025).
@@
Table 6: The latency and the speedup of different linear layers in PixArt-$\alpha$/$\Sigma$ under different quantization precision generating a $1024\times 1024$ image. Result measured on RTX 5080 GPU using CUDA 12.8.
@@ 5 Conclusion
In this paper, we propose FP4DiT, a PTQ method that achieves W4A6 and W4A8 quantization on T2I DiT using FPQ. We use a mixed FP formats strategy based on the special structure of DiT and propose scale-aware AdaRound to enhance the weight quantization for FPQ. We analyze the difference between the activation of U-Net DM and DiT and apply token-wise online activation quantization based on the findings. Our experiments demonstrate the superior performance of FP4DiT compared to other quantization methods on the quantative HPSv2 benchmark, MS-COCO dataset and qualitative visualization comparison at minimial hardware cost.
@@ A.2 Additional Ablation Studies
1:Full-precision array $A_{\text{FP}}$, number of bits $n$, exponent bits $n_{e}$, mantissa bits $n_{m}$, clipping value $maxval$
2:$A_{\text{abs}}\leftarrow\text{abs}(A_{\text{FP}})$
3:$bias\leftarrow 2^{n_{e}}-\log_{2}(A_{\text{abs}})+\log_{2}(2-2^{-n_{m}})-1$
4:$A_{\text{clip}}\leftarrow\min(\max(A_{\text{FP}},-maxval),maxval)$
5:$S_{\text{log}}\leftarrow\text{clamp}(\lfloor\log_{2}(\text{abs}(A_{\text{clip}}))\rfloor+bias,1)$
6:$S\leftarrow 2.0^{(S_{\text{log}}-n_{m}-bias)}$
7:$result\leftarrow\text{round-to-nearest}(A_{\text{clip}}/S)\times S$
8:return $result$
@@ A.3.1 Floating Point Quantization Scheme
Since Floating Point Quantization (FPQ) is not as straightforward as INT quantization, there has not been a simple and unified algorithm for performing FPQ yet. In this paper, we apply Algorithm 1 from (Kuzmin et al., 2022) to perform the FPQ. Unlike (Kuzmin et al., 2022) learning the clipping value $maxval$ and bit allocations between mantissa and exponent part, we use the absolute maximum of the tensor as $maxval$ and perform FPQ with a predetermined FP format.
@@ A.4 Evaluation Settings
For CLIP score (Hessel et al., 2021), we apply the openai-clip  [^1] https://github.com/openai/CLIP library to measure the CLIP score between prompts and generated images. We employ ViT-B/32 as the CLIP model. To measure HPSv2 (Wu et al., 2023a) score, we generate 3.2k images across the four categories (800 images per category) and use the provided human preference predictor to estimate the performance. For the Fréchet Inception Distance (FID) (Heusel et al., 2017) measurement, we apply clean-fid  [^2] https://github.com/GaParmar/clean-fid library to measure the FID between generated images and ground-truth images.
@@ A.6 Additional Visualization Results
In this section, we present the random samples derived from full-precision and W4A6/W4A8 PixArt-$\alpha$, PixArt-$\Sigma$, and Hunyuan that are quantized by different baseline and FP4DiT. As depicted by Figures 11, 12, 13, 14, 15, and 16, FP4DiT generates results of impressive visual content. FP4DiT consistently shows superior performance across various DiT models. We list a detailed comparison between FP4DiT and other baselines as follows:
@@
    PixArt-$\alpha$: On W4A8 (Fig. 11), FP4DiT generates more fine-grained images than all other baselines, such as the texture of macarons and waves. Besides, other baselines fail to generate a hedgehog while FP4DiT correctly depicts it. On W4A6 (Fig. 12), our method shows near W4A8 performance, while other baselines become noisy and lose details.
@@
    Hunyuan: On W4A8 (Fig. 15), FP4DiT’s generated images more closely align with the full precision model. For example, in the first image, FP4DiT has the same color right face while Q-Diffusion and TMFQ-DM alter the face’s color. In the third image, FP4DiT generates the most similar face decoration. In addition, FP4DiT also has the least noise on Hunyuan DiT. On W4A6 (Fig. 16), all methods generate blur images. However, the contour of images identified from FP4DiT’s visualization results matches the full precision images, while other baselines fail to produce a contour or the contour is not prompt-adherent.

### v3

@@ 4.2 Main Results
In our experiment, we consider four baseline approaches: Q-Diffusion (Li et al., 2023), TFMQ-DM (Huang et al., 2024), ViDiT-Q (Zhao et al., 2024) and Q-DiT (Chen et al., 2025a). Note that while Q-Diffusion and TFMQ-DM are originally designed for U-Nets, we modify these approaches to use the same online token-wise activation quantization as FP4DiT per Table 1, while ViDiT-Q and Q-DiT uses this mechanism by default. Further, note that ViDiT-Q uses mix-precision meaning some of their layers are not quantized to 4-bits. For the sake of convenience, we use their mix-precision as a W4 baseline to conduct our experiments.
@@
We generate $512\times 512$ resolution images using PixArt-$\alpha$ and $1024\times 1024$ for PixArt-$\Sigma$ and Hunyuan. For evaluation, we primarily consider the Human Preference Score v2 (HPSv2) (Wu et al., 2023a) benchmark and MS-COCO 2014 (Lin et al., 2014a) validation set. Specifically, HPSv2 considers four image categories: animation, concept-art, painting and photography, and estimates the human preference score of an image generated using a prompt with respect to one category. Each category contains 800 prompts, requiring 3.2k images to be generated to fully evaluate. The final HPSv2 score is the average estimated human preference across all four categories. Additionally, for COCO 2014, we measure the FID (Heusel et al., 2017), CLIP (Hessel et al., 2021) score and average ImageReward (IR) (Xu et al., 2023) score using prompts from the validation set.
@@
Table 2 compares our method with the stated baseline approaches at the W4A8 and W4A6 precision levels. FP4DiT consistently outperforms all other methods across three different base models at every precision level in terms of each performance metric. On HPSv2, FP4DiT achieves higher performance across every model and bit-precision combination. Although other approaches may achieve slightly higher performance on an individual category, these gaps are small, while FP4DiT leads in terms of FID, and CLIP on the COCO 2014 validation set. Finally, we observe a modest amount of IR degradation for PixArt-$\alpha$ and Hunyuan compared to the W16A16 model, however this is most pronounced for PixArt-$\Sigma$ where all quantization methods suffer a substantial IR score loss.
We believe this degradation is two-fold. First, PixArt-$\Sigma$ generates larger images, which means there is more pixel space to introduce noise. Second, IR estimates how much a human would prefer an image given the prompt, as opposed to CLIP which measures prompt adherence by comparing the latent embeddings of prompt and image, meaning the errors introduced by quantization are likely to have a greater impact. Nevertheless, FP4DiT achieves the highest IR performance in most of the model and precision settings.
@@
Next, we compare performance across activation bit-precision. For the PixArt-series DiT models, A6 precision can cause substantial performance degradation for some baselines, while FP4DiT still maintains high performance. This is specifically noteworthy on PixArt-$\Sigma$ as baseline approaches suffer a loss in terms of either FID and/or CLIP performance at A6 compared to A8, but this is not the case for FP4DiT. For the Hunyuan DiT model, A6 quantization remains highly challenging across the board, however, our method still achives the best results at this level except the IR score. Overall though, Table 2 demonstrates the efficacy of FP4DiT in term of quantitative T2I compared to prominant baseline approaches.
@@
Next, Figure 7 provides qualitative image samples on the PixArt models. Note the higher quality of the images generated by FP4DiT at both the A8 and A6 levels. Specifically, on PixArt-$\alpha$, the puppies have more realistic detail and the generated image more closely aligns with the W16A16 model. This is especially true for the PixArt-$\Sigma$ sample images, where FP4DiT accurately captures the intricate details of the portrait of a middle-aged Asian woman, including the fractured porcelain and paint splatters. In contrast, the other quantization methods either introduce significant artifacts, lose fine details, or fail to accurately represent the prompt’s content, further highlighting the robustness and superior performance of the FP4DiT approach.
@@ 4.4 Hardware Cost Comparison
Table 6: The latency and the speedup of different linear layers in PixArt-$\alpha$$\Sigma$ under different quantization precision generating a $1024\times 1024$ image, e.g., batch size of 2, 4096 tokens, each with an embedding dimension of 1152. Result measured on RTX 5080 GPU using CUDA 12.8.
@@
We develop a CUDA 12.8 kernel and measure hardware latency on an Nvidia 5080, a Blackwell GPU which supports low-bit FP formats. We consider the self-attention weight/activation dimensions found in PixArt-$\alpha$$\Sigma$ when generating an $1024\times 1024$ image. Table 6 reports our findings. The result indicates that, despite the general expectation that floating-point computations are more demanding than integer operations, the latency results are nearly identical across all tested quantization precisions. Specifically, although A6 precision imposes some additional overhead compared to A8 (either for FP or INT), this is quite minor, leading to an overall speedup around 1.5x compared to W16A16, much like the popular W4A8 integer quantization (Li et al., 2023; Zhao et al., 2024; Chen et al., 2025a).
@@ 5 Conclusion
In this paper, we propose FP4DiT, a PTQ method that achieves W4A6 and W4A8 quantization on T2I DiT using FPQ. We use a mixed FP formats strategy based on the special structure of DiT and propose scale-aware AdaRound to enhance the weight quantization for FPQ. We analyze the difference between the activation of U-Net DM and DiT and apply token-wise online activation quantization based on the findings. Our experiments demonstrate that FP4DiT achieves performance improvement in terms of quantative benchmarks such as CLIP and ImageReward scores on the MS-COCO dataset and HPSv2 benchmark, while also obtaining qualitative visualization improvement at minimal hardware cost.
@@ A.2 Additional Ablation Studies
1:Full-precision array $A_{\text{FP}}$, number of bits $n$, exponent bits $n_{e}$, mantissa bits $n_{m}$, clipping value maxvalmaxval
[…]
8:return resultresult
A.2.2 Effect of Calibration Dataset
N=64 N=64 N=64 N=64
25.09 27.43 28.21 28.20
Table 10: The quantization results for different sizes of calibration dataset with PixArt-$\alpha$.
We investigate FP4DiT’s effectiveness under different calibration dataset sizes $N$. Table 10 reports the HPSv2 score of the W4A8 PixArt-$\alpha$ quantized by our method with $N=64,128,256,$ and $512$. The results show steady improvements as $N$ increases, with performance saturating beyond $N=256$. In particular, FP4DiT achieves a strong score of 27.43 with only 128 samples, and further gains to 28.21 at $N=256$, while larger calibration does not bring additional benefits. This demonstrates that FP4DiT requires only a relatively small calibration set to achieve near-optimal performance, making it both efficient and effective compared to existing approaches.
A.2.3 Classifier-Free Guidance (CFG)
Precision CFG=3.0 CFG=4.5 CFG=7.0
W4A8 27.22 27.43 26.74
Full Precision 30.97 31.01 31.26
Table 11: The quantization results for different CFG scales with PixArt-$\alpha$.
We investigate FP4DiT’s effectiveness under different CFG scales. Table 11 compares FP4DiT’s quantization performance with PixArt-$\alpha$ for three different CFG scales, which are 3.0, 4.5 (default), and 7.0. Although FP4DiT is calibrated at $\text{CFG}=4.5$, it remains robust and exhibits only minor degradation under different CFG scales, which is still outstanding among the baseline methods as shown in Table 2.
@@ A.3.1 Floating Point Quantization Scheme
Since Floating Point Quantization (FPQ) is not as straightforward as INT quantization, there has not been a simple and unified algorithm for performing FPQ yet. In this paper, we apply Algorithm 1 from (Kuzmin et al., 2022) to perform the FPQ. Unlike (Kuzmin et al., 2022) learning the clipping value maxvalmaxval and bit allocations between mantissa and exponent part, we use the absolute maximum of the tensor as maxvalmaxval and perform FPQ with a predetermined FP format.
@@ A.4 Evaluation Settings
For CLIP score (Hessel et al., 2021), we apply the openai-clip  [^1] https://github.com/openai/CLIP library to measure the CLIP score between prompts and generated images. We employ ViT-B/32 as the CLIP model. To measure HPSv2 (Wu et al., 2023a) score, we generate 3.2k images across the four categories (800 images per category) and use the provided human preference predictor to estimate the performance. For the Fréchet Inception Distance (FID) (Heusel et al., 2017) measurement, we apply clean-fid  [^2] https://github.com/GaParmar/clean-fid library to measure the FID.
@@ A.6 Additional Visualization Results
In this section, we present the random samples derived from full-precision and W4A6/W4A8 $\alpha, PixArt$-$\Sigma$, and Hunyuan that are quantized by baselines and FP4DiT. As depicted by Figures 11, 12, 13, 14, 15, and 16, FP4DiT generates results of impressive visual content. FP4DiT consistently shows superior performance across various DiT models. We list a detailed comparison between FP4DiT and the baselines:
@@
    PixArt-$\alpha$: On W4A8 (Fig. 11), FP4DiT generates more fine-grained images, e.g., macaroon and wave details. Besides, the baselines fail to generate a hedgehog while FP4DiT correctly depicts it. On W4A6 (Fig. 12), our method shows near W4A8 performance, unlike the noisier baselines.
@@
    Hunyuan: On W4A8 (Fig. 15), FP4DiT’s generated images closely align with the full precision model. For example, in the first image, FP4DiT has the same color right face while Q-Diffusion and TMFQ-DM alter the face’s color. In the third image, FP4DiT generates the most similar face decoration. In addition, FP4DiT also has the least noise on Hunyuan DiT. On W4A6 (Fig. 16), all methods generate blur images. However, the contour of images identified from FP4DiT’s visualization results matches the full precision images, while other baselines fail to produce a prompt-adherent contour.